# Hybrid optimal control with mixed-integer Lagrangian methods

### v1

@@ 8 $\varepsilon_{j+1}\leftarrow\max\{\epsilon_{\textup{d}},\kappa_{\varepsilon}\varepsilon_{j}\}$
Algorithm 3.1 Augmented Lagrangian routine for (P) using [5, Algorithm 3.1] as inner solver
The major computational toll is taken by Algorithm 3.1, which aims to minimize $\mathcal{L}_{\mu_{j}}(\cdot,\widehat{y}^{j})$ over $\mathcal{X}$. In fact, only approximate criticality is required, and the sequential mixed-integer linearization scheme of [5] can be readily adopted to compute a suitable critical point $x^{j}$. Although the performance of the algorithm might be improved by warm starting the inner solver and guaranteeing descent behaviour [5], these properties are not strictly required (convergence results are typically not affected). Algorithms 3.1 and 3.1 incorporate a classical first-order dual estimate update. The dual safeguard takes place in Algorithm 3.1 where the compact set $\mathcal{Y}$ can be a generic hyperbox or tailored to the constraint set $\mathcal{C}$ to exploit additional known structures [6, 24].
Given a (possibly inexact, first-order) solution $x$ to (5), the dual update rule in Algorithm 3.1 is designed toward the identity
@@
as usual in AL methods. This relationship allows to monitor the “outer” convergence for (P) with the “inner” subproblem tolerance for (5). Since the inclusion $y\in\mathcal{N}_{\mathcal{C}}(s)$ in Definition 2.4 is always satisfied by construction of $s^{j}$ and $y^{j}$ in Algorithms 3.1 and 3.1, approximate KKT criticality can be directly detected in Algorithm 3.1 returning a pair $(x^{j},y^{j})$.
Finally, penalty parameter and inner tolerance are adapted in Algorithms 3.1 and 3.1 following classical update rules [2, 6]. Initial values for steering primal-dual tolerance sequences can be user-specified or adaptively selected based on the initial infeasibility and criticality measure [2, 8]. We shall mention also that, when a feasible initial point is available, reset schemes are applicable and provide asymptotic feasibility guarantees [9] hardly attainable otherwise.
In analogy to magical steps [2, 4] and feasibility pumps [3], additional steps can be optionally included in Algorithm 3.1 to ease or accelerate the solution process. Such routines could pursue feasibility (switching to the minimization of an infeasibility measure) or refine the incumbent solution by solving a nonlinear program with fixed integer-valued variables [5].
@@ 4 Numerical Results
In this section, we present results achieved by the proposed numerical approach for several high-dimensional hybrid optimal control problems. Before going deeper into details, it is important to highlight some aspects regarding the computational procedure behind Algorithm 3.1.
For starting the AL scheme, we set $y^{0}=0$ and $\mu_{1}=\varepsilon_{1}=10^{-2}$ for the initial multiplier, penalty parameter and inner tolerance, respectively. The choice of $x^{0}$ is based on a more sophisticated approach. Admissible points for the set $\mathcal{X}$ are computed as projections onto it, in the $\ell_{1}$-norm sense. Another routine was implemented to help generating reasonable starting points and refining solutions provided by Algorithm 3.1. While integer variables are relaxed in the former case, they are fixed (to an admissible value) in the latter case so that only real-valued variables are further optimized. In both scenarios, one ends up with a nonlinear problem, which is then tackled by Ipopt. We use Gurobi for solving MIL subproblems and the numerical routine of [5] for executing Algorithm 3.1.
@@ 4.1 Car with hysteretic turbo charger and drag
The first problem handled by the proposed method is inspired by [18, Section V.A] and [5, Section 4.1]. The aim is to perform point-to-point one-dimensional motion planning of a car with hybrid nonlinear dynamics extending previous models to accommodate a drag force term. The underlying double-integrator point-mass model is equipped with a hysteretic turbo accelerator. The car is described by its time-dependent position $x(t)$, velocity $v(t)$ and a turbo state $w(t)\in\{0,1\}$ for each time point $t\in[0,T]$, $T>0$. The control variables correspond to the inputs to the acceleration $a(t)$ and brake $b(t)$ pedals. The turbo mode becomes active or inactive once the velocity exceeds $v^{+}:=10$ or falls below the threshold $v^{-}:=5$, respectively. An activated turbo mode entails an effective increase of the nominal thrust by three times but leaves the braking force unaffected. These dynamics read
$\dot{x}(t)={}$ $v(t),$
$\dot{v}(t)={}$ $\tau(w(t),a(t))-b(t)-c_{D}v(t)^{2}$
where $c_{D}\geq 0$ denotes the drag coefficient, and the thrust $\tau$ has two modes of operation depending on the turbo state: $\tau(w,a)=a$ if $w=0$, and $\tau(w,a)=3a$ if $w=1$.
State and control bounds, boundary conditions, the fixed time interval and the minimum-control objective functions are chosen as in [5, §4.1]. The same holds for the hysteresis characteristic describing the turbo behaviour whose logical implications are encoded as MIL constraints based on big-M reformulations.
@@
Considering the final time $T=10$, the problem is fully discretized using the explicit Euler scheme on a uniform time grid with $N=90$ intervals. The resulting problem in the form (P) has $637$ optimization variables (of which $91$ integer-valued) and $91$ nonlinear constraints.
Figure 1 shows the time-dependent position and velocity for different values of the drag coefficient $c_{D}\in[0,0.003]$. In all tests, boundary conditions and path constraints are satisfied. For $c_{D}=0$, the results obtained coincide with those presented in [5] for the linear-only case. However, setting the drag coefficient to the largest value in the range, i.e. $c_{D}=0.003$, prevents the car from reaching its maximum velocity $v_{\max}=25$.
To report on the numerical performance of the method, we consider the cases with $c_{D}=0$ and $c_{D}>0$ separately. In the first (all linear) case, Algorithm 3.1 terminated after 12 (outer) iterations and 11 (wall clock) seconds requiring 338 (cumulative) calls to the MIL solver. For $c_{D}>0$, the problem becomes nonlinear, and the respective values are in the range of 42–44 iterations, 21–36 seconds, and 965–1710 MIL calls.
Figure 1: Position $x(t)$ (left) and velocity $v(t)$ (right) of a car with a hysteretic turbo charger and drag force. The turbo mode is also displayed as active (dots) or inactive (crosses).
@@ 4.2 Lotka-Volterra fishing with total variation
In this subsection, we present results for a Lotka-Volterra fishing problem with a total variation (TV) and a fixed final time. For two different populations $x_{1}$ and $x_{2}$, which can be interpreted as the biomasses of prey and predator, respectively, binary optimal control is performed. The primary objective is formulated as
$\min\quad\int_{0}^{T}\|x(t)-x_{r}\|^{2}\mathrm{d}t$ (9)
for tracking the reference value $x_{r}=\begin{bmatrix}1,1\end{bmatrix}^{\top}$ [23, §5]. The dynamics of the underlying 2-population model can be described by the equations
@@
for given parameters $c_{1},c_{2}\in\mathbb{R}^{5}$ and controls $w(t)\in\{0,1\}^{5}$ [22]. The pointwise constraint $\sum_{i=1}^{5}w_{i}(t)=1$ encodes the choice of one among five fishing options.
To limit the well-known chattering behaviour of the optimal binary controls, we monitor the total variation (in fact, the discrete counterpart thereof)
$TV(w)\approx\frac{1}{2}\sum_{i=1}^{m}\sum_{k=0}^{N-1}|w_{i,k+1}-w_{i,k}|,$ (10)
possibly imposing an upper bound $TV(w)\leq U_{TV}$ or including a penalty term $\alpha_{TV}TV(w)$ in the objective [23].
@@
The initial state is chosen as $x(0)=\begin{bmatrix}0.5,0.7\end{bmatrix}^{\top}$. The mixed-integer OCP is fully discretized using the explicit Euler scheme over a uniform time grid with final time $T=12$ and $N=30$ intervals. The objective’s integral is approximated by the left Riemann sum. The resulting discretized problem contains $398$ optimization variables (of which $155$ integer-valued) and $31$ nonlinear constraints.
The initial guess values are set to $1$ for all variables. In order to obtain a warm start, the problem (with no TV cost nor limit) is considered with relaxed integrality and solved by Ipopt. Such solution, displayed in Figure 2, is then projected onto $\mathcal{X}$ and used as input $x^{0}$ for Algorithm 3.1 (for all subsequent tests).
Figure 2: Initial guess for the fishing problem, before projection.
We now consider two problems to highlight the effect of penalizing and limiting the TV: with TV cost $\alpha_{TV}=1$, we first solve an instance with no TV upper bound ($U_{TV}=\infty$) and then with $U_{TV}=10$. After executing Algorithm 3.1, its solution is refined with Ipopt by optimizing over real-valued variables. The corresponding results are shown in Figures 3 and 4, respectively. Limiting the (discretized) TV helps to avoid a chattering behaviour of the optimal solution without significantly impairing the tracking performance. Indeed, the solution with the TV constraint appears to improve both components of the objective, namely tracking and switching. This phenomenon indicates that the results in Figure 3 correspond to a suboptimal solution testifying the local nature of Algorithm 3.1.
Figure 3: Refined solution for the fishing problem with TV cost $\alpha_{TV}=1$ and no TV limit ($U_{TV}=\infty$).
Figure 4: Refined results for a fishing problem with TV cost $\alpha_{TV}=1$ and TV limit $U_{TV}=10$.
As to the numerical performance of Algorithm 3.1, the MIL solver was called 6764 times for the case without TV limit returning a solution after 42 iterations and 150 seconds. For $U_{TV}=10$, these statistics read 2869 MIL calls, 41 iterations and 61 seconds.
4.3 Dynamic facility location problem
Let us now consider a dynamic facility location problem with applications in logistics. Consider a set of $m$ customers and $n$ possible warehouses. Each customer is given by a position $x_{j}\in\mathbb{R}^{2}$ and demand $d_{j}$, $j\in\{1,\ldots,m\}$. Each potential warehouse is associated to a region $R_{i}\subset\mathbb{R}^{2}$, its position $x_{i}$, capacity $u_{i}$, and a binary $b_{i}$ encoding whether it is actually built or not, $i\in\{1,\dots,n\}$. The amount of goods delivered from warehouse $i$ to customer $j$ (to satisfy their demand) is denoted by $l_{ij}$ whose location-dependent transportation cost is denoted by $C_{ij}(x_{i})$. Moreover, building a warehouse $i$ incurs the capacity-dependent construction cost $P_{i}(u_{i})$.
The aim is to determine, which potential warehouses are to be built as well as their position and capacity such that the total cost is minimized:
$\operatorname*{minimize}~\sum_{i=1}^{n}b_{i}\left(P_{i}(u_{i})+\sum_{j=1}^{m}C_{ij}(x_{i})\right)$ (11)
where $P_{i}$ and $C_{ij}$ are given construction and transportation cost functions, for $i\in\{1,\ldots,n\}$, $j\in\{1,\ldots,m\}$. Several constraints are incorporated into the model. First, urban planning restrictions have to be taken into account: $x_{i}\in R_{i}$ for each warehouse. Next, capacity specifications need to be integrated: $\underline{u}_{i}b_{i}\leq u_{i}\leq\overline{u}_{i}b_{i}$ and $\sum_{j=1}^{m}l_{ij}\leq u_{i}$, with $0\leq\underline{u}_{i}\leq\overline{u}_{i}$ the capacity limits for warehouse $i$. Moreover, customer demands have to be satisfied: $\sum_{i=1}^{n}l_{ij}=d_{j}$ for each customer. Lastly, integrality and nonnegativity conditions are required: $b_{i}\in\{0,1\}$ and $l_{ij}\geq 0$, respectively, for each warehouse $i$ and customer $j$. In order to decouple integer-valued variables and nonlinear functions, these relationships are rewritten by introducing some auxiliary variables and using big-M reformulations.
Simulations
We present results for the problem scenario with $n=30$ regions and $m=10$ customers, depicted in Figure 5, with $1080$ optimization variables (of which $30$ integer-valued) and $330$ nonlinear constraints. Regions $R_{i}$ are placed along two circles and are described by simple bounds for $x_{i}$. Regions are numbered from $1$ to $n$ starting with those at the inner circle and moving counter-clockwise. The capacity bounds are set to $\underline{u}_{i}=10$ and $\overline{u}_{i}=1000$. The construction cost function is chosen as $P_{i}(u):=iu+c_{P}$, $i\in\{1,\ldots,n\}$ where $c_{P}$ is a fixed nonnegative offset. As to the transportation cost function, we consider $C_{ij}(x_{i}):=c_{C}\sqrt{\operatorname{dist}(x_{i},x_{j})}$ where $c_{C}:=(4\sqrt{3\sqrt{2\alpha_{u}}})/3$, $\alpha_{u}=0.1$. Such cost corresponds to the analytical solution for a point-to-point motion (between $x_{i}$ and $x_{j}$) with free final time minimizing both final time and (integral squared) acceleration with cost factors $1$ and $\alpha_{u}$, respectively.
Figure 5 shows a feasible initial guess and the solution obtained for $c_{P}=100$, which selects 7 warehouses that overall strike a good balance between construction and transportation costs. For this instance, Algorithm 3.1 terminated after 42 iterations and 67 seconds calling the MIL solver 936 times.
Finally, we considered instances with increasing values for $c_{P}$. The optimal number of warehouses depending on parameter $c_{P}$ is shown in Figure 6 and, despite some oscillations due to local minima, the overall behaviour is consistent with our expectations in that the number of warehouses decreases with increasing $c_{P}$.
Figure 5: Facility location problem with $n=30$ regions for warehouses (gray rectangles), $m=10$ customers (red triangles), and $c_{P}=100$: position of warehouses actually built (black dots) in initial guess (left) and solution obtained (right).
Figure 6: Facility location problem with $n=30$ regions for warehouses and $m=10$ customers: number of warehouses built as a function of $c_{P}$.

### v2

@@ 8       $\varepsilon_{j+1}\leftarrow\max\{\epsilon_{\textup{d}},\kappa_{\varepsilon}\varepsilon_{j}\}$
Algorithm 3.1 Safeguarded augmented Lagrangian routine for (P)
The major computational toll is taken by Algorithm 3.1, which aims to minimize $\mathcal{L}_{\mu_{j}}(\cdot,\widehat{y}^{j})$ over $\mathcal{X}$. In fact, only approximate criticality is required, and the sequential mixed-integer linearization scheme of [6, Algorithm 3.1] can be readily adopted to compute a suitable $\varepsilon_{j}$-critical point $x^{j}$, in the sense of Definition 2.2. Although the performance of the algorithm might be improved by warm-starting the inner solver and guaranteeing descent behaviour, these properties are not strictly required in the convergence analysis [4]. The classical first-order dual estimate update takes place at Algorithms 3.1 and 3.1. Given a (possibly inexact, first-order) solution $x$ to (5), the dual update rule in Algorithm 3.1 is designed toward the identity
@@
as usual in AL methods. This relationship allows to monitor the “outer” convergence for (P) with the “inner” subproblem tolerance for (5). Since the inclusion $y\in\mathcal{N}_{\mathcal{C}}(s)$ in Definition 2.4 is always satisfied by construction of $s^{j}$ and $y^{j}$ in Algorithms 3.1 and 3.1, approximate KKT criticality can be directly detected in Algorithm 3.1 returning a pair $(x^{j},y^{j})$. The dual safeguard takes place in Algorithm 3.1, where the compact set $\mathcal{Y}$ can be a generic hyperbox or tailored to the constraint set $\mathcal{C}$ to exploit additional known structures. Finally, penalty parameter and inner tolerance are adapted in Algorithms 3.1 and 3.1 following classical update rules [1, §12.4]. Initial values for steering primal-dual tolerance sequences can be user-specified or adaptively selected based on the initial infeasibility and criticality measure. We shall mention also that, when a feasible initial point is available, reset schemes are applicable and provide asymptotic feasibility guarantees hardly attainable otherwise.
In analogy to magical steps [1, §8.2], [5, §4.1] and feasibility pumps [3], additional steps can be optionally included in Algorithm 3.1 to ease or accelerate the solution process. Such routines could pursue feasibility (switching to the minimization of an infeasibility measure) or refine the incumbent solution by solving a nonlinear program with fixed integer-valued variables [6, Rem. 3.1], [13, §5].
Remark 3.1.
A favorable attribute of AL methods is that if the AL subproblems are solved to approximate global optimality, then the iterates accumulate at feasible, globally optimal points [1, Thm 5.2 ]. Thus, whenever additional knowledge or problem structure allow to solve the AL subproblems to approximate global optimality, this general property of AL methods automatically provides global guarantees for (P).
@@ 4 Numerical Results
In this section, we present and compare the results achieved by the proposed numerical approach for two hybrid optimal control problems. Before considering the simulations, we highlight some aspects regarding the computational procedure behind Algorithm 3.1 (ALM).
ALM is initialized with zero dual estimate $y^{0}$, penalty parameter $\mu_{1}=10^{-2}$ and inner tolerance $\varepsilon_{1}=10^{-2}$. The choice of $x^{0}$ is based on a more sophisticated approach. Admissible points for the set $\mathcal{X}$ are computed as projections onto it, in the $\ell_{1}$-norm sense. Another routine was implemented to generate reasonable starting points and to refine solutions provided by ALM. While integer variables are relaxed in the former case, they are fixed (to an admissible value) in the latter case so that only real-valued variables are further optimized. In both scenarios, one ends up with a nonlinear problem, which is then tackled by Ipopt [19]. We use the numerical routine of [6] for executing Algorithm 3.1, with Gurobi for solving the MIL subproblems there.
@@ 4.1 Car with hysteretic turbo charger and drag
The first problem handled by the proposed method is inspired by [14, §V.A] and [6, §4.1]. The aim is to perform point-to-point one-dimensional motion planning of a car with hybrid nonlinear dynamics, extending previous models to accommodate a drag force term. The underlying double-integrator point-mass model is equipped with a hysteretic turbo accelerator. The car is described by its time-dependent position $x(t)$, velocity $v(t)$ and a turbo state $w(t)\in\{0,1\}$ for each time point $t\in[0,T]$, $T>0$. The control variables correspond to the inputs to the acceleration $a(t)$ and brake $b(t)$ pedals. The turbo mode becomes active or inactive once the velocity exceeds $v^{+}:=10$ or falls below the threshold $v^{-}:=5$, respectively. An activated turbo mode entails an effective increase of the nominal thrust by three times but leaves the braking force unaffected. These dynamics read
$\dot{x}(t)={}$ $v(t),$ $\dot{v}(t)={}$ $\tau(w(t),a(t))-b(t)-c_{\rm d}v(t)^{2}$
where $c_{\rm d}\geq 0$ denotes the drag coefficient, and the thrust $\tau$ has two modes of operation depending on the turbo state: $\tau(w,a)=a$ if $w=0$, and $\tau(w,a)=3a$ if $w=1$. The velocity is limited by $|v(t)|\leq v_{\max}:=25$. Boundary conditions, state and control bounds, and the minimum-control objective function are chosen following [6, §4.1]. The same holds for the hysteresis characteristic describing the turbo behaviour, whose logical implications are encoded as MIL constraints.
@@
Considering the final time $T=10$, the problem is fully discretized using the explicit Euler scheme on a uniform time grid with $N=100$ intervals. The resulting problem in the form (P) has $100$ binary-valued optimization variables and $101$ nonlinear constraints.
ALM is initialized with an all-zero primal input. The solutions returned for different values of the drag coefficient $c_{\rm d}\in\{10^{-3},10^{-2}\}$ are displayed in Figure 1, together with the solution obtained with dynamic programming (DP). For DP, we used uniform discretization grids with 100 values for each (real-valued) state and control. In all tests, boundary conditions and path constraints are approximately satisfied. As expected, all solutions resemble the one presented in [6] for the linear case with $c_{\rm d}=0$, whereas more acceleration and less braking are demanded with stronger drag. The DP solution exhibits some chattering behaviour in the braking phase—this result may be an artifact due to discretization. However, finer discretizations appeared impractical in terms of runtime (on a standard laptop): for each call, DP took minutes to compute, whereas ALM was in the order of seconds.
For $c_{\rm d}=10^{-3}$, the turbo car problem was in addition solved by ALM combined with a subsequent fixed-integer NLP refinement to examine the impact of the problem size on the computational effort. As illustrated in Figure 2, the NLP runtime grows faster but is lower than ALM’s, which however increases only linearly with $N$.
Figure 1: Solutions for the turbo car problem with different values of drag coefficient $c_{\rm d}\in\{10^{-3},10^{-2}\}$. Comparison of Algorithm 3.1 (ALM), starting from an all-zero initial guess, and dynamic programming (DP).
Figure 2: Runtimes for the turbo car problem with drag $c_{\rm d}=10^{-3}$ for different number $N$ of discretization intervals. Solutions obtained with ALM and subsequent fixed-integer NLP refinement.
@@ 4.2 Lotka-Volterra fishing with total variation
Let us now consider the classical Lotka-Volterra fishing problem complemented with a total variation ($\operatorname{TV}$) term [17]. For two states $x_{1}$ and $x_{2}$, which can be interpreted as the biomasses of prey and predator populations, respectively, a binary optimal control is sought. The dynamics are described with the model
@@
for some given parameters $c_{1},c_{2}\in\mathbb{R}^{5}$ and control inputs $w(t)\in\{0,1\}^{5}$. For every time $t\in\mathbb{R}$, the pointwise constraint $\sum_{i=1}^{5}w_{i}(t)=1$ encodes the choice of one among five control options. The primary objective is formulated as
$J(x):=\min\quad\int_{0}^{T}\|x(t)-x_{r}\|^{2}\mathrm{d}t$
for tracking the reference value $x_{r}=(1,1)$, starting from the initial state $x_{1}(0)=0.5$, $x_{2}(0)=0.7$ and with final time $T=12$ [18, §5].
To limit the well-known chattering behaviour of the optimal control, we monitor the total variation (in fact, its discrete counterpart, after time discretization)
$\operatorname{TV}(w):=\frac{1}{2}\sum_{i=1}^{5}\sum_{k=1}^{N-1}|w_{i,k+1}-w_{i,k}|,$
possibly imposing an upper bound $\operatorname{TV}(w)\leq U_{\operatorname{TV}}$ or including a penalty term $\alpha_{\operatorname{TV}}\operatorname{TV}(w)$ in the objective. By including additional $(N-1)$ real-valued variables and $4(N-1)$ linear inequality constraints for each binary-valued control, the nonsmooth $\operatorname{TV}$ term can be expressed as a linear function of the unknowns.
@@
The mixed-integer OCP is fully discretized using the explicit Euler scheme over a uniform time grid with $N=50$ intervals. The resulting discretized problem contains $250$ integer-valued optimization variables and $51$ nonlinear constraints.
Figure 3: Solutions for the fishing problem without $\operatorname{TV}$, starting CIA and ALM from the NLP relaxation.
The initial guess values are set to $1$ for all variables and the problem (with no TV cost nor limit) is first considered with relaxed integrality and solved by Ipopt. Such solution is then used as input for the combinatorial integral approximation (CIA) [18], computed via pycombina [2], and for warm-starting ALM. The refined solutions are displayed in Figure 3 and summarized in Table 1, together with the globally optimal solution obtained with DP (on a uniform grid with 100 values for each state). We report the effective control $\Omega:=\sum_{i=1}^{5}iw_{i}$, instead of the vector-valued $w$, to better display the selected fishing option. When warm-started from the NLP relaxation, the infeasible initial guess is first projected onto $\mathcal{X}$ before running ALM; the returned solution corresponds to such projection and does not provide good tracking performance. Similarly, when warm-started from the DP or CIA solution, ALM returns the same control sequence, validating their local optimality.
Table 1: Results for the fishing problem with and without $\operatorname{TV}$: upper bound $U_{\operatorname{TV}}$ on the total variation $\operatorname{TV}$, tracking cost $J$. When not indicated, ALM starts from the CIA solution.
solver $U_{\operatorname{TV}}$ $\operatorname{TV}(w)$ $J(x)$
DP $\infty$ 14 0.388
ALM from DP $\infty$ 14 0.388
NLP relaxation $\infty$ 2.6 0.381
ALM from NLP $\infty$ 3 0.834
CIA from NLP $\infty$ 38 0.398
ALM $\infty$ 38 0.398
solver $U_{\operatorname{TV}}$ $\operatorname{TV}(w)$ $J(x)$
ALM 38 38 0.398
ALM 33 33 0.415
ALM 28 28 0.825
ALM 23 23 0.549
ALM 18 18 1.206
ALM 13 13 0.601
Figure 4: Solutions for the fishing problem with limited TV, starting ALM from the CIA solution.
We now consider the performance of ALM when limiting the control $\operatorname{TV}$. Always starting from the CIA solution, ALM is executed with decreasing values $U_{\operatorname{TV}}$ of maximum admissible $\operatorname{TV}$. The corresponding results are shown in Figure 4 and collected in Table 1. Note that since the ALM solution coincides with the CIA solution for $U_{\operatorname{TV}}=38$, this case is omitted in the plots from Figure 4.
Expectedly, limiting the (discretized) $\operatorname{TV}$ helps avoiding a chattering behaviour of the optimal solution but at the cost of impairing the tracking performance. In all cases, the returned solution had the maximum TV allowed. Moreover, Table 1 reports that the solution for $U_{\operatorname{TV}}=23$ (resp. $U_{\operatorname{TV}}=13$) had a better tracking objective than the one for $U_{\operatorname{TV}}=28$ ($U_{\operatorname{TV}}=18$), testifying the local nature of ALM when coupled with the local inner solver of [6].